# Agile Tradespace Exploration for Space Rendezvous Mission Design via Transformers

### v1

@@ Abstract
Spacecraft rendezvous enables on-orbit servicing, debris removal, and crewed docking, forming the foundation for a scalable space economy. Designing such missions requires rapid exploration of the tradespace between control cost and flight time across multiple candidate targets. However, multi-objective optimization in this setting is challenging, as the underlying constraints are often highly nonconvex, and mission designers must balance accuracy (e.g., solving the full problem) with efficiency (e.g., convex relaxations), slowing iteration and limiting design agility. To address these challenges, this paper proposes an AI-powered framework that enables agile mission design for a wide range of Earth orbit rendezvous scenarios. Given the orbital information of the target spacecraft, boundary conditions, and a range of flight times, this work proposes a Transformer-based architecture that generates, in a single parallelized inference step, a set of near-Pareto optimal trajectories across varying flight times, thereby enabling rapid mission trade studies. The model is further extended to accommodate variable flight times and perturbed orbital dynamics, supporting realistic multi-objective trade-offs. Validation on chance-constrained rendezvous problems with passive safety constraints demonstrates that the model generalizes across both flight times and dynamics, consistently providing high-quality initial guesses that converge to superior solutions in fewer iterations. Moreover, the framework efficiently approximates the Pareto front, achieving runtimes comparable to convex relaxation by exploiting parallelized inference. Together, these results position the proposed framework as a practical surrogate for nonconvex trajectory generation and mark an important step toward AI-driven trajectory design for accelerating preliminary mission planning in real-world rendezvous applications. [^1] The code will be released upon the camera-ready version.
@@ 1 Introduction
Figure 1: This paper presents a framework for AI-powered agile mission design. Mission configurations (target orbits, boundary conditions, and flight-time ranges) serve as inputs to a trajectory generation pipeline (top), where a Transformer model autoregressively predicts control sequences that are propagated through the system dynamics (middle). Batch inference efficiently produces multiple Pareto-optimal trajectories across flight times that are encoded into a tensor of [B(batch size),N(time horizon),D(subsequence length)][B($\text${batch size}),N($\text${time horizon}),D($\text${subsequence length})], which can be projected onto the multi-objective solution space as surrogate mission models. Selected point solutions can then be further refined using sequential convex programming (bottom).
@@
Historically, space mission design has been an arduous, highly manual process heavily dependent on expert knowledge. This difficulty arises from two main factors: (i) its formulation as a multi-objective optimization problem, in which several competing objectives (e.g., time of flight, fuel consumption, radiation exposure, observability) must be optimized simultaneously, and (ii) the presence of complex nonconvex and often implicit constraints, such as safety and observation requirements. The result is a constrained Multi-Objective Optimal Control Problem (MO-OCP), where the Pareto front of feasible solutions is of central importance. Because general MO-OCPs are extremely challenging to solve, traditional practice has relied on problem-specific search strategies [8] or metaheuristics [9], typically under strong simplifying assumptions, to produce preliminary trajectories of acceptable quality. However, while useful in early mission design, these methods suffer from limited scalability and generalizability. Moreover, they require a full recomputation whenever mission parameters change, underscoring their fragility in dynamic design environments.
@@
To address these limitations, this study introduces an agile Transformer-based trajectory generation pipeline for RPOD mission design (cf. Figure 1). This framework is capable of generating batches of near-Pareto-optimal and near-feasible trajectories in a single inference pass. This approach builds on the Autonomous Rendezvous Transformer (ART) framework [15][16][17][18], originally developed as a warm-starting technique for nonconvex trajectory optimization to accelerate onboard guidance. In this work, ART is extended and repurposed for mission design. Specifically, for conducting trade studies in complex scenarios and for approximating the Pareto front of inherently nonconvex RPOD optimization problems.
@@
  • This paper proposes a framework for AI-powered—namely, Transformer-based—agile space mission design that generates a near-Pareto optimal solution set for a given mission configuration.
  • This paper extends the ART architecture to meet the requirements of mission design. Specifically, this work investigates (i) the generalization across dynamics and flight time, and (ii) the generation of Pareto-optimal trajectories via parallelizable inference.
  • Numerical experiments validate these requirements through an extensive set of rendezvous mission profiles in perturbed-Keplerian orbits, incorporating data from the latest space debris catalog. Crucially, trajectory generation with ART achieves runtimes comparable to those of convex relaxation methods, underscoring its effectiveness as a practical and agile tool for mission design.
The remainder of this paper is organized as follows. Section 2 reviews prior work on mission design and learning-based trajectory generation, while Section 3 introduces the necessary background on ART. Section 4 presents the formulation of the MO-OCP and introduces the ART extension for Pareto-optimal trajectory generation. Section 5 formulates the spacecraft rendezvous problem in perturbed Keplerian orbits as a chance-constrained nonconvex optimization problem. Section 6 presents the generalization performance of the proposed framework and analyzes the resulting multi-objective solution surface. Finally, Sec. 7 concludes the paper and discusses directions for future work.
@@ 2.1 Space Mission Design
These include deterministic broad search under problem-specific simplifying assumptions (e.g., patched-conics with impulsive maneuvers [8][19]), tree-based motion planning techniques [20][21], and stochastic, population-based methods such as genetic algorithms and differential evolution [9][22][23][24][25]. While capable of exploring global design spaces and accommodating uncertainty in high-fidelity simulations [26][27][28], these methods generally require extensive tuning, suffer from poor scalability, and provide limited generalization.
@@
Another strategy is to reduce the multi-objective problem to a scalarized one, either through weighted sums or more sophisticated scalarization schemes [29]. Such formulations can then be solved via conventional nonconvex methods [30][31], often leveraging algorithms such as Sequential Convex Programming (SCP) [32]. Hybrid methods combining stochastic exploration with gradient-based local refinement have also been explored [29]. However, approximating the Pareto set typically requires sweeping large sets of weights or thresholds, resulting in high computational cost. Moreover, incorporating expressive constraints, such as signal temporal logic or mixed-integer conditions [33][34][35], further increases the risk of non-convergence.
@@
Simplified dynamics and surrogate models are often employed to accelerate the design of DSS [4][36]. While these methods reduce computational burden, they may neglect operational constraints (e.g., passive safety [10][37][38][39], field-of-view/approach cones [40][41][42], plume impingement effects [33], etc.) that strongly influence feasibility. As a result, preliminary designs generated under these simplifications often require significant refinement. Crucially, these methods typically do not yield full trajectories but only approximate mission profiles.
In summary, there remains a gap between computationally intensive, high-fidelity trajectory design tools and the fast but coarse approximations provided by surrogate models. This paper seeks to bridge that gap by introducing an AI–based trajectory generation framework that accelerates the design process while preserving generality, without requiring re-training when mission configurations change.
@@ 2.2 Learning-based Trajectory Generation in Space
In recent years, there has been a growing interest in applying machine learning and AI techniques to generate spacecraft trajectories for both mission design and onboard autonomy contexts. Learning-based methods have been applied in different capacities, including the prediction of optimal co-states for indirect methods [43], identifying active constraint sets [44], or estimating primal-dual variables [45]. Direct trajectory generation has also been pursued via imitation learning [46] and reinforcement learning [47][48], leveraging both Transformer- [15] and diffusion-based architectures [49][50]. Neural architectures have been applied to highly multi-modal trajectory optimization problems, such as Earth–Moon low-thrust transfers, in both direct [51] and indirect formulations [52], including cases with missed-thrust contingencies [53]. Many of these methods are used to provide warm-starts [54][55] for nonlinear optimization problems, which have recently been demonstrated on Astrobee aboard the International Space Station [56]. Within this context, ART has extended Transformer-based trajectory generation to a variety of settings, including model predictive control [16], chance-constrained fault-tolerant optimal control [18], and trajectory optimization across varying flight time and dynamic obstacle scenarios [17]. Despite these advances, existing learning-based frameworks still lack two critical capabilities for practical mission design: (i) robust generalization across orbital regimes, and (ii) the ability to generate and explore Pareto-optimal trade-offs in multi-objective settings. This paper addresses these limitations by developing a framework for generalizable nonconvex trajectory generation that approximates the true Pareto set.
@@ 3 Background: Autonomous Rendezvous Transformer
ART [15] is a trajectory generation framework that leverages the sequence modeling and prediction capabilities of Transformer models. Given a dataset of state and control trajectories, the core idea is to train the model via next-element prediction, learning to generate the optimal control inputs and corresponding states at each step. A key enabler of ART for capable trajectory generation is the state representation—that is, the tokenization of continuous trajectory information into sequences of states, controls, and associated performance metrics. In ART, the tokenized trajectory representation is typically defined as:
@@
where $\boldsymbol{g}_{k}$ denotes the goal state, and $\boldsymbol{x}_{k}$ and $\boldsymbol{u}_{k}$ represent the state and control input at time step $k$, respectively. $N$ indicates the length of the time horizon, which has a constant time displacement $\Delta t$ for each time step. Additionally, $\boldsymbol{o}_{k}$ represents a scenario-specific observation vector at time step $k$ (e.g., information about obstacles in the scene). This work considers two specific performance metrics, $\{r_{k},c_{k}\}$, where $r_{k}\in\mathbb{R}$ is the Reward-To-Go (RTG) and $c_{k}\in\mathbb{Z}_{\geq 0}$ is the Constraint-To-Go (CTG). These metrics are defined as follows:
@@
where $j_{j}$ is the instantaneous reward and $\boldsymbol{1}_{\mathcal{S}_{j}}(\cdot)$ is the indicator function for the feasible domain at the time step $k$. Essentially, both RTG and CTG measure the cumulative future optimality and feasibility of the trajectory starting from the current state. A subsequence is defined as a tuple $\boldsymbol{g}_{k},\boldsymbol{x}_{k},{r}_{k},{c}_{k},\boldsymbol{o}_{k},\boldsymbol{u}_{k}]$, where each element type is embedded into a common input space of dimension $D=6$, corresponding to the number of distinct element types. These embeddings are then projected into the Transformer’s latent space. This general framework allows tokenization of trajectory optimization problems with an arbitrary number of objectives and constraints.
@@
Dataset Generation: The process begins by generating a set of randomized scenario configurations through domain randomization of key problem parameters (e.g., boundary conditions, flight times). For each configuration, a corresponding trajectory optimization problem is solved to obtain reference solutions for training. Diversity across parameters, operating conditions, and performance metrics is essential, as it ensures exposure to a broad solution landscape and supports generalization of the underlying dynamics and control strategies.
Training: After collecting a sufficient number of trajectories, the Transformer model is trained to minimize a supervised loss. The final layer of the Transformer branches into two task-specific heads, each implemented as a linear layer: one for state prediction and one for control input prediction. The loss function over a batch of $B$ trajectories with horizon $N$ is defined as:
@@ 4.1 Multi-Objective Optimal Control Problem
Space mission design involves identifying feasible solutions while optimizing multiple, often competing, mission objectives. In the context of RPOD missions, a fundamental trade-off in preliminary trajectory design is balancing fuel consumption against flight time. Given this consideration, it is crucial for mission designers to characterize the Pareto front of the associated Multi-Objective Optimal Control Problem (MO-OCP):
@@
There are two primary challenges in solving the MO‑OCP above. First, multi-objective optimization problems are typically solved with computationally intensive metaheuristics [26][27]. A common alternative is to reformulate the problem as a single-objective optimization using either a weighted sum of objectives or a constraint-based scalarization [23], where all but one objective are converted into an inequality constraint with upper bounds. However, these reformulated problems can yield nondominated solution sets that are highly sensitive to minor variations in constraints, thereby limiting their robustness and generalizability. Moreover, the problem is often nonconvex due to the presence of nonlinear dynamics and nonconvex feasible sets for the state and control variables. SCP offers a promising solution strategy by iteratively solving convex approximations of the nonconvex problem. This method leverages efficient convex solvers but necessitates a sufficiently accurate initial guess to ensure rapid convergence and to avoid poor local minima, as highlighted by Malyuta et al. [32].
@@ 4.2 ART for Efficient Pareto-Optimal Trajectory Generation
Since ART was originally developed for single-objective, single-constraint problems, several extensions are introduced to enable trajectory generation over a multi-objective solution space. Specifically, the dimensions of the RTG and CTG are generalized to accommodate arbitrary numbers of rewards and constraints, such that $\boldsymbol{r}_{k}\in\mathbb{R}^{n_{r}}$ and $\boldsymbol{c}_{k}\in\mathbb{Z}_{\geq 0}^{n_{c}}$ denote the vectorized RTG and CTG, respectively. To achieve generalization across flight times and environmental parameters, the approach proposed by Celestini et al. [17] is adopted. Dataset generation employs domain randomization over three elements: (i) boundary conditions, (ii) flight times, and (iii) observation parameters defining the system dynamics. In particular, flight time is controlled by a fixed step size $\Delta t$ together with a variable horizon length $N$, ensuring that the Transformer’s learned dynamics evolve under a consistent transition interval. This formulation simplifies training and improves efficiency at inference by allowing variable-length trajectories to be stacked and autoregressively generated within a single forward pass (cf. Figure 1). Specifically, variable-horizon trajectories $\boldsymbol{\tau}_{1:N}^{(1:B)}$ are organized into a tensor of shape $[B,N,D]$, where $N$ denotes the maximum horizon length in the batch. Shorter trajectories are padded to match the same length. At each step, the model autoregressively produces state–control input tokens, yielding $[B,D]$ outputs that are assembled into complete trajectories.

### v2

@@ Abstract
Spacecraft rendezvous enables on-orbit servicing, debris removal, and crewed docking, forming the foundation for a scalable space economy. Designing such missions requires rapid exploration of the tradespace between control cost and flight time across multiple candidate targets. However, multi-objective optimization in this setting is challenging, as the underlying constraints are often nonconvex, and mission designers must balance accuracy (e.g., solving the full problem) with efficiency (e.g., convex relaxations), slowing iteration and limiting design agility. To address these challenges, this paper proposes an AI-powered framework that enables agile and generalized rendezvous mission design. Given the orbital information of the target spacecraft, boundary conditions of the servicer, and a range of flight times, a transformer model generates a set of near-Pareto optimal trajectories across varying flight times in a single parallelized inference step, thereby enabling rapid mission trade studies. The model is further extended to accommodate variable flight times and perturbed orbital dynamics, supporting realistic multi-objective trade-offs. Validation on chance-constrained rendezvous problems in Earth orbits with passive safety constraints demonstrates that the model generalizes across both flight times and dynamics, consistently providing high-quality initial guesses that converge to superior solutions in fewer iterations. Moreover, the framework efficiently approximates the Pareto front, achieving runtimes comparable to convex relaxation by exploiting parallelized inference. Together, these results position the proposed framework as a practical surrogate for nonconvex trajectory generation and mark an important step toward AI-driven trajectory design for accelerating preliminary mission planning in real-world rendezvous applications.
@@ 1 Introduction
Figure 1: This paper presents a framework for AI-powered agile mission design. Mission configurations (target orbits, boundary conditions, and flight-time ranges) serve as inputs to a trajectory generation pipeline (top), where a transformer model autoregressively predicts control sequences that are propagated through the system dynamics (middle). Batch inference efficiently produces multiple Pareto-optimal trajectories across flight times that are encoded into a tensor of [B(batch size),N(time horizon),D(types of subsequence elements)][B($\text${batch size}),N($\text${time horizon}),D($\text${types of subsequence elements})], which can be projected onto the multi-objective solution space as surrogate mission models. Selected point solutions can then be further refined using sequential convex programming (bottom).
@@
Historically, space mission design has been an arduous, highly manual process, largely dependent on expert knowledge and rarely regarded as a component of onboard autonomy. This difficulty arises from two main factors: (i) its formulation as a multi-objective optimization problem, in which several competing objectives (e.g., flight time, fuel consumption, radiation exposure, observability) must be optimized simultaneously, and (ii) the presence of complex nonconvex and often implicit constraints, such as safety and observation requirements. The result is a constrained Multi-Objective Optimal Control Problem (MO-OCP), where the Pareto front of feasible solutions is of central importance. Because general MO-OCPs are extremely challenging to solve, traditional practice has relied on problem-specific search strategies [34] or metaheuristics [38], typically under strong simplifying assumptions, to produce preliminary trajectories of acceptable quality. However, while useful in early mission design, these methods suffer from limited scalability and generalizability. Moreover, they require a full recomputation whenever mission parameters change, underscoring their fragility in dynamic design environments.
@@
To address these limitations, this study introduces an agile transformer-based trajectory generation pipeline for RPOD mission design (cf. Figure 1). This framework is capable of generating batches of near-Pareto-optimal and near-feasible trajectories in a single inference pass. This approach builds on the Autonomous Rendezvous Transformer (ART) framework [23][12][11][57][55], originally developed as a warm-starting technique for nonconvex trajectory optimization to accelerate onboard guidance. In this work, ART is extended and repurposed for mission design, specifically for conducting trade studies in complex scenarios and for approximating the Pareto front of nonconvex RPOD trajectory optimization problems.
@@
  • This paper proposes a framework for AI-powered (transformer-based) agile space mission design that generates a near-Pareto optimal solution set for a given mission configuration.
  • This paper extends the ART architecture to meet the requirements of space mission design. Specifically, this work investigates (i) the generalization across dynamics and flight time, and (ii) the generation of Pareto-optimal trajectories via parallelizable inference.
  • Numerical experiments validate these requirements through an extensive set of rendezvous mission profiles in perturbed-Keplerian orbits, incorporating data from the latest space debris catalog. Crucially, trajectory generation with ART achieves runtimes comparable to those of convex relaxation methods, underscoring its effectiveness as a practical and agile tool for multi-objective nonconvex trajectory generation.
The remainder of this paper is organized as follows. Section 2 reviews prior work on mission design and learning-based trajectory generation, while Section 3 introduces the necessary background on ART. Section 4 presents the formulation of the MO-OCP and introduces the ART extension for Pareto-optimal trajectory generation. Section 5 formulates the spacecraft rendezvous problem in perturbed Keplerian orbits as a chance-constrained nonconvex optimization problem. Section 6 presents the generalization performance of the proposed framework and analyzes the resulting multi-objective solution surface. Finally, Section 7 concludes the paper and discusses directions for future work.
@@ 2.1 Space Mission Design
One of the popular approaches is the deterministic broad search under problem-specific simplifying assumptions (e.g., patched-conics with impulsive maneuvers [34][59], tree-based motion planning techniques [52][15]). While state-of-the-art algorithms enable a fast broad search, crude approximation leads to a non-negligible gap between the approximated objective function during the initial guess search and the true objective function in the later refinement [54]. Another common approach is using stochastic, population-based methods such as genetic algorithms and differential evolution [38][29][62][17][30]. While capable of exploring global design spaces and accommodating uncertainty in high-fidelity simulations [2][58][19], these methods generally require extensive tuning, suffer from poor scalability, and provide limited generalization.
@@
Another strategy is to reduce the multi-objective problem to a scalarized one, either through weighted sums or more sophisticated scalarization schemes [63]. Such formulations can then be solved via nonconvex optimization methods [5][45], often leveraging algorithms such as Sequential Convex Programming (SCP) [41]. Hybrid methods combining stochastic exploration with gradient-based local refinement have also been explored [63]. However, approximating the Pareto set typically requires sweeping large sets of weights or thresholds, resulting in high computational cost. Moreover, incorporating expressive constraints, such as signal temporal logic or mixed-integer constraints [42][53][43], further increases the risk of non-convergence.
@@
Simplified dynamics and surrogate models are often employed to accelerate the design of DSS [40][39]. While these methods reduce computational burden, they may neglect operational constraints (e.g., passive safety [13][1][22][60], field-of-view/approach cones [31][35][56], plume impingement effects [42], etc.) that strongly influence feasibility. As a result, preliminary designs generated under these simplifications often require large corrections. Crucially, these methods typically do not yield full trajectories but only approximate mission profiles.
In summary, there remains a gap between computationally intensive, high-fidelity trajectory design methods and the fast but coarse approximations provided by surrogate models. This paper seeks to bridge that gap by introducing an AI–based trajectory generation framework that accelerates the design process while preserving generality. The proposed approach is highly flexible and capable of predicting accurate Pareto solution surfaces under complex constraints, without requiring re-training the neural network model when mission configurations change.
@@ 2.2 Learning-based Trajectory Generation in Space
In recent years, there has been a growing interest in applying machine learning and AI techniques to generate spacecraft trajectories for both mission design and onboard autonomy contexts. This is primarily motivated by the goal of facilitating trajectory generation by (i) reducing the overall computational cost, and (ii) enabling global search over candidate trajectories, which conventional methods often struggle to achieve without extensive expert hand-tuning. Learning-based methods have been applied in different capacities, including the prediction of optimal co-states for indirect methods [28], identifying active constraint sets [9], or estimating primal-dual variables [7]. Direct trajectory generation has also been pursued via imitation learning [61] and reinforcement learning [10][66], leveraging both transformer- [23] and diffusion-based architectures [8][37]. Neural architectures have been applied to highly multi-modal trajectory optimization problems, such as Earth–Moon low-thrust transfers [36] [20], including cases with missed-thrust contingencies [51]. Many of these methods are used to provide warm-starts [4][65] for nonlinear optimization problems, which have recently been demonstrated on Astrobee aboard the International Space Station [3]. Within this context, ART has extended transformer-based trajectory generation to a variety of settings, including model predictive control [12], chance-constrained and fault-tolerant optimal control [57], and trajectory optimization across varying flight time and dynamic obstacle scenarios [11]. More recently, nonconvex trajectory generation based on natural language inputs with trajectory-level safety is demonstrated by the authors [55]. Despite these advances, existing learning-based frameworks still lack two critical capabilities for practical space mission design: (i) robust generalization across orbital regimes, and (ii) the ability to generate and explore Pareto-optimal trade-offs in multi-objective settings. This paper addresses these limitations by developing a framework for generalizable nonconvex trajectory generation that approximates the true Pareto set.
@@ 3 Background: Autonomous Rendezvous Transformer
ART [23] is a trajectory generation framework that leverages the sequence modeling and prediction capabilities of causal transformer models. Given a dataset of state and control trajectories, the core idea is to train the model via next-element prediction, learning to generate the optimal control inputs and corresponding states at each step. A key enabler of ART for capable trajectory generation is the state representation, which is the tokenization of continuous trajectory information into sequences of states, controls, and associated performance metrics. In ART, the tokenized trajectory representation is typically defined as
@@
where $\boldsymbol{g}_{k}$ denotes the goal state, and $\boldsymbol{x}_{k}$ and $\boldsymbol{u}_{k}$ represent the state and control input at time step $k$, respectively. $N$ indicates the length of the time horizon, which has a constant time displacement $\Delta t$ for each time step. Additionally, $\boldsymbol{o}_{k}$ represents a scenario-specific observation vector at time step $k$ (e.g., information about obstacles in the scene). This work considers two specific performance metrics, $\{r_{k},c_{k}\}$, where $r_{k}\in\mathbb{R}$ is the Reward-To-Go (RTG) and $c_{k}\in\mathbb{Z}_{\geq 0}$ is the Constraint-To-Go (CTG). These metrics are defined as
@@
where $j_{j}$ is the instantaneous reward and $\boldsymbol{1}_{\mathcal{S}_{j}}(\cdot)$ is the indicator function for the feasible domain at the time step $j$. Essentially, both RTG and CTG measure the cumulative future optimality and feasibility of the trajectory starting from the current state. A subsequence is defined as a tuple $\boldsymbol{g}_{k},\boldsymbol{x}_{k},{r}_{k},{c}_{k},\boldsymbol{o}_{k},\boldsymbol{u}_{k}]$, where each element type is embedded into a common input space of dimension $D=6$, corresponding to the number of distinct element types. These embeddings are then projected into the transformer’s latent space. This general framework allows tokenization of trajectory optimization problems with an arbitrary number of objectives and constraints.
@@
Dataset Generation: The process begins by generating a set of scenario configurations through domain randomization of key problem parameters (e.g., boundary conditions, flight times). For each configuration, a corresponding trajectory optimization problem is solved to obtain reference solutions for training. Diversity across parameters, operating conditions, and performance metrics is essential, as it ensures exposure to a broad solution landscape and supports generalization of the underlying dynamics and control strategies.
Training: After collecting a sufficient number of trajectories, the causal transformer model is trained to minimize a supervised loss. The final layer of the transformer branches into two task-specific heads, each implemented as a linear layer: one for state prediction and one for control input prediction. The loss function over a batch of $B$ trajectories with horizon $N$ is defined as
@@ 4.1 Multi-Objective Optimal Control Problem
Space mission design involves identifying feasible solutions while optimizing multiple, often competing, mission objectives. In the context of RPOD missions, a fundamental trade-off in preliminary trajectory design is balancing fuel consumption against flight time. Given this consideration, it is crucial for mission designers to characterize the Pareto front of the associated MO-OCP:
@@
There are two primary challenges in solving the above MO-OCP. First, multi-objective optimization problems are typically solved with computationally intensive metaheuristics [2][58]. A common alternative is to reformulate the problem as a single-objective optimization using either a weighted sum of objectives or a constraint-based scalarization [62], where all but one objective are converted into an inequality constraint with upper bounds. However, these reformulated problems can yield nondominated solution sets that are highly sensitive to minor variations in constraints, thereby limiting their robustness and generalizability. Moreover, the problem is often nonconvex due to the presence of nonlinear dynamics and nonconvex feasible sets for the state and control variables. SCP offers a promising solution strategy by iteratively solving convex approximations of the nonconvex problem. This method leverages efficient convex solvers but necessitates a sufficiently accurate initial guess to ensure rapid convergence and to avoid poor local minima, as highlighted by Malyuta et al. [41].
@@ 4.2 ART for Efficient Pareto-Optimal Trajectory Generation
Since ART was originally developed for single-objective, single-constraint problems, several extensions are introduced to enable trajectory generation over a multi-objective solution space. Specifically, the dimensions of the RTG and CTG are generalized to accommodate arbitrary numbers of rewards and constraints, such that $\boldsymbol{r}_{k}\in\mathbb{R}^{n_{r}}$ and $\boldsymbol{c}_{k}\in\mathbb{Z}_{\geq 0}^{n_{c}}$ denote the vectorized RTG and CTG, respectively. To achieve generalization across flight times and environmental parameters, the approach proposed by Celestini et al. [11] is adopted. Dataset generation employs domain randomization over three elements: (i) boundary conditions, (ii) flight times, and (iii) observation parameters defining the system dynamics. The randomization domain directly determines the domain over which the model generalizes, as induced by the seen data during the training. In particular, flight time is controlled by a fixed step size $\Delta t$ together with a variable horizon length $N$, ensuring that the transformer’s learned dynamics evolve under a consistent transition interval. This formulation simplifies training and improves efficiency at inference by allowing variable-length trajectories to be stacked and autoregressively generated within a single forward pass (cf. Figure 1). Specifically, variable-horizon trajectories $\boldsymbol{\tau}_{1:N}^{(1:B)}$ are organized into a tensor of shape $[B,N,D]$, where $N$ denotes the maximum horizon length in the batch. Shorter trajectories are padded to match the same length. At each step, the model autoregressively produces state–control input tokens, yielding $[B,D]$ outputs that are assembled into complete trajectories.